# GenMix: Effective Data Augmentation with Generative Diffusion Model Image Editing

### v3

@@ Abstract
Data Augmentation, Mixup Data Augmentation, Diffusion Models, Generative Synthetic Data, Synthetic Data
1 Introduction
@@
Deep Neural Networks (DNNs) have achieved remarkable advancements across various computer vision tasks, such as image classification Dosovitskiy et al. (2020); He et al. (2016a); Liu et al. (2022); Tan and Le (2019), object detection Fang et al. (2021); He et al. (2017); Wang et al. (2021), image captioning Fei et al. (2022); Ke et al. (2019), human pose estimation Liu et al. (2021); Luo et al. (2021), and image segmentation Badrinarayanan et al. (2017); Strudel et al. (2021). However, as the complexity of these tasks increases, so do the computational costs associated with modern DNN architectures, including larger memory footprints, model sizes, higher FLOPs, and longer inference times DeVries and Taylor (2017). Recent models employ billions of parameters to enhance feature representation. This often leads to overfitting and challenges in generalization, especially when the training data is limited or poorly structured. These issues become even more pronounced in cross-domain settings, where models must adapt to divergent data distributions.
@@
Image mixing based data augmentation techniques have proven effective for enhancing the generalization of deep learning models. A wide array of such techniques has been introduced, including CutMix Yun et al. (2019a), Mixup Kim et al. (2020a), SnapMix Huang et al. (2021), Co-Mixup Kim et al. (2020b), SaliencyMix Uddin et al. (2020), GuidedMixup Kang and Kim (2023), MixPro Zhao et al. (2022), TransMix Chen et al. (2022a), and PuzzleMix Kim et al. (2020a). These methods typically involve combining portions of randomly selected images of corresponding labels through various mixing strategies to generate augmented samples - see top row in Fig. 1. By employing linear interpolation or other blending techniques, they create novel training images to enrich the dataset for addressing potential model overfitting Yun et al. (2019b); Zhang et al. (2018); Han et al. (2022); Qin et al. (2023); Huang et al. (2021); Islam (2022).
@@
Though effective, image mixing based data augmentation encounters challenges, such as the omission of salient regions and label ambiguities due to the random blending of image content Kim et al. (2020a). To mitigate these issues, several approaches have incorporated saliency-driven methods, which prioritize the preservation of important regions by overlaying them onto less significant background areas of the target image Kim et al. (2020a); Uddin et al. (2020); Kang and Kim (2023); Islam et al. (2022). While these techniques offer improvements, they introduce considerable computational overhead and rely on the accuracy of the saliency methods, which is a limiting factor. Furthermore, by combining images from diverse categories, these techniques can still overlook crucial contextual information, leaving the problem of inadequate feature preservation unresolved.
@@
Figure 1: Representative examples of augmentation techniques. Top: Conventional image mixing methods like CutMix Yun et al. (2019b), Mixup Zhang et al. (2018), GridMix Baek et al. (2021), SmoothMix Lee et al. (2020), ResizeMix Qin et al. (2020) and FMix Harris et al. (2020) use interpolation or other strategies to mix source and target images, largely disregarding the natural appearance of the resulting image. Bottom: Our GMix is a label-preserving technique like AugMix Hendrycks et al. (2019), PixMix Hendrycks et al. (2022) and IPMix Huang et al. (2023). However, it utilizes a generative diffusion model with custom tailored prompts for high quality augmentation. Three different outputs (1A, 2A and 3A) for the target are shown for GenMix.
@@
Diffusion Models (DMs) Takagi and Nishimoto (2023); Du et al. (2023); Luo et al. (2023); Saharia et al. (2022); Dhariwal and Nichol (2021) have recently emerged as powerful tools for image-to-image generation and editing. Some studies Trabucco et al. (2024); Azizi et al. (2023) have also explored using DM generated images to augment training data. However, we empirically observe limited performance gains from these approaches, with models trained on synthetic data occasionally underperforming compared to the baseline models without augmentation Azizi et al. (2023). This can be attributed to the limited control over image generation of the DMs. Due to the model sensitivity to conditional prompts, generating complex scenes, layouts, and shapes becomes challenging Zhang et al. (2023). Poorly crafted prompts often yield unsuitable images for augmentation, which can lead to model adaption to an incorrect data distribution. This underscores the need for more precise prompt selection and better control mechanisms in the augmentation process to improve both in-domain generalization and cross-domain performance.
@@
To address this, we propose GenMix, a generlizable data augmentation technique that leverages generative pre-trained diffusion models to compute diverse augmentation samples through carefully designed conditional prompts. GMix creates hybrid images by blending original and generated content, maintaining essential semantics of the original image, see bottom row of Fig. 1. To further enrich structural diversity, we incorporate self-similarity fractals into the hybrid images, a strategy proven to enhance machine learning safety Hendrycks et al. (2022); Huang et al. (2024). This blending mitigates overfitting and improves model performance. Our strategy is unlike any previous attempts of directly using DM images for augmentation Azizi et al. (2023); Trabucco et al. (2024). It effectively incorporates a range of operations in its components to achieve outstanding eventual performance - see Table I. Our experiments show that GMix surpasses current state-of-the-art (SOTA) augmentation methods by improving generalization and adversarial robustness. Additionally, it facilitates prompt-guided learning in both in-domain and cross-domain tasks, making it highly adaptable for diverse computer vision problems. Furthermore, our results confirm that GMix is compatible with a wide range of datasets and it integrates seamlessly into existing SOTA frameworks.
@@
  • We explore the applicability of GMix to both in-domain classification and domain adaptation tasks, introducing sets of bespoke conditional prompt for enhanced adaptability. Notably, GMix is the first method to address both in-domain and domain adaptation tasks effectively.
  • We demonstrate with extensive experiments across eight datasets that GMix outperforms existing SOTA augmentation methods in general classification, fine-grained classification, adversarial robustness, and cross-domain tasks.
@@
Diff-Mix
Aziz
DA-Fusion
@@
Components Source image
@@
Tasks Adversarial robustness
@@
A preliminary version of this research was presented at IEEE CVPR 2024 Islam et al. (2024). This article is a major extension of the preliminary work. Here, we propose a more effective approach to fuse the original and generated images by introducing a smoothness factor, leading to consistent performance gains. As compared to Islam et al. (2024), we also extend the fractal image dataset to add more variety and structural complexity to the the final augmented images. Moreover, besides providing considerable further details for the technique, prior art and evaluation, we also explore the applicability of our work to domain adaption tasks such as open-partial, open-set, and partial-set to facilitate a more comprehensive view of the resulting data augmentation. To the best of our knowledge, GenMix is the first augmentation technique that enables significant performance gains in both in-domain and cross-domain tasks, signifying its importance in training robust large-scale learning systems.
2 Related Work
@@
Diffusion Models for Augmentation. Recently, there has been a growing interest in leveraging Diffusion Models (DMs) for data augmentation. Azizi et al. Azizi et al. (2023) proposed using text-to-image fine-tuned diffusion models to generate synthetic samples for ImageNet classification, demonstrating that incorporating these generated images into the training set can enhance classification performance. Similarly, Trabucco et al. Trabucco et al. (2024) utilized off-the-shelf diffusion models to create diverse and semantically rich images via prompt engineering, aiming to improve image classification on various datasets. Furthermore, Li et al. Li et al. (2023) explored using diffusion models for data augmentation in conjunction with knowledge distillation, particularly in scenarios where real images are unavailable. These efforts underscore the potential of diffusion models to enrich data augmentation techniques.
Image Mixing for Augmentation. Image mixing has become a widely employed category of data augmentation techniques, enhancing the performance and robustness of deep learning models Liang et al. (2023); Yan et al. (2023); Mensink and Mettes (2023); Chen and Lu (2023). Notable state-of-the-art methods in this category include CutMix, AugMix and PuzzleMix. CutMix Yun et al. (2019a) improves out-of-distribution generalization by randomly overlaying a patch from a source image onto a target image. AugMix Hendrycks et al. (2019) stochastically combines augmentation operations like equalize, posterize, and translate-x to introduce semantic diversity. PuzzleMix Kim et al. (2020a) refines mixup techniques by considering image saliency and local statistics. Similarly, another technique SaliencyMix Uddin et al. (2020) leverages saliency maps to focus on the most important image regions, preserving their integrity. Manifold Mixup Verma et al. (2019) blends hidden network states during training, creating an interpolated state for improved feature representation. Another well-known method is PixMix Hendrycks et al. (2022) that extends AugMix Hendrycks et al. (2019) by mixing input images with fractal and feature visualization images, improving machine learning safety. A comprehensive overview of various image-mixing methods, along with their components and task accomplishment is presented in Table I.
Automated Data Augmentation. Several researchers have also explored automated data augmentation for improving model performance. For instance, AutoAugment Cubuk et al. (2018) utilizes reinforcement learning to discover optimal data augmentation strategies, while RandAugment Cubuk et al. (2020) employs a set of randomized augmentation operations to enhance model generalization and robustness. AdaAug Cheung and Yeung (2021) has been introduced to efficiently learn adaptive augmentation policies that vary depending on the class or even the instance.
@@
Figure 2: Overview of the proposed GMix method. An input image and a randomly selected prompt are input to a pre-trained image editing diffusion model to obtain a generated image. Input and generated images are smoothly concatenated using a binary mask to obtain a hybrid image. A random fractal image is finally blended with this hybrid image to obtain the augmented image.
3 Proposed Method
3.1 Overview
Existing image mixing based data augmentation techniques may induce label ambiguities by pasting the source image contents onto the target image and consequently replacing either important regions of the foreground or the required background context Yun et al. (2019b); Kang and Kim (2023). The central idea of our approach is to augment the source image only while preserving the key semantics and concatenating a portion of the source image with its generative counterpart. At the same time, our method introduces diverse object details and contexts in the resulting image for effective augmentation.
The schematics of the proposed GMix technique is provided in Fig. 2. Our method comprises three main steps: prompt-guided image editing, seamless concatenation, and fractal interpolation. Initially, we use conditional prompts with a diffusion-based image editing module to compute generative counterparts of the source image. Then, a portion of the source image is concatenated with the rest of the generated image, forming a hybrid image. This approach ensures that the training model has access to the original visual content along with the generated one. Finally, the hybrid image is blended with a random fractal image, obtaining the final training image with a diverse structure. Improvements of our image by blending fractal is inspired by Hendrycks et al. (2022, 2021). However, in our case, the blending objective for augmentation is unique to our approach.
3.2 GenMix Details
The proposed GMix is quite effective data augmentation method that aims to enhance the generalization and robustness of the deep neural networks. Formally, $I_{i}\in\mathbb{R}^{h\times w\times c}$ is a training dataset image. Our data augmentation method is defined as: $\mathcal{D}$mix$(\cdot):\mathbb{R}^{h\times w\times c}\to\mathbb{R}^{h\times w\times c}$. Input image $I_{i}$ goes through proposed generation using randomly selected prompt $p_{j}$, seamless concatenation using random mask $M_{u}$, and linearly interpolated using a random fractal image $F_{v}$ to obtain the final augmented image $A_{ijuv}$. The overall augmentation process, as shown in Algorithm 1, may be represented as
@@ 3.2 GenMix Details
Prompt Guided Image Editing: Our image editing step $\mathcal{E}(.)$ consists of a pre-trained image editing diffusion model that requires a prompt $p_{j}$ from a predefined set of $k$ prompts, $P=\{p_{1},p_{2},\dots,p_{k}\}$ where $j\in[1,k]$ along with the input training image $I_{i}$ and produces an edited counterpart image $\hat{I}_{ij}$. The image generation process in conventional diffusion models is often open-ended for image-to-image or text-to-image translation. On the other hand, image editing is guided by textual prompts. The aim is to obtain diverse image-to-image translations guided by textual prompts. In our case, as the goal is to achieve a slightly modified but not too different version of $I_{i}$, filter-like prompts are curated in $\mathcal{P}$ which do not alter the image drastically. Examples of the prompts used in GMix are shown in Figure 3. The overall generation step can be represented as:
@@
$A_{ijuv}=\lambda F_{v}+(1-\lambda)H_{iju},\text{ }0\leq\lambda<1$ (4)
where $\lambda$ is the interpolation factor. This results in the final augmented image $A_{ijuv}$ used to train or fine-tune deep neural networks. The overall augmentation process of GMix can be represented as:
Algorithm 1 GMix
1:$I_{i}\in\mathcal{D}$ training images dataset, $m$: number of augmented images, $p_{j}\in\mathcal{P}$ set of textual prompts, $M_{u}\in\mathcal{M}$ set of smooth masks, $F_{v}\in\mathcal{F}$ set of self-similarity fractal images, $\lambda$: interpolation weight
2:$\mathcal{D}^{\prime}$: $m$ Augmented images
3:$\mathcal{D}^{\prime}\leftarrow\emptyset$

### v4

@@ Abstract
Data Augmentation, Diffusion Models, Generative Models, Synthetic Data, Image Editing
I Introduction
@@
Deep Neural Networks (DNNs) have achieved remarkable advancements across various computer vision tasks, such as image classification [1, 2, 3, 4], object detection [5, 6, 7], image captioning [8, 9], human pose estimation [10, 11], and image segmentation [12, 13]. However, as the complexity of these tasks increases, so do the computational costs associated with modern DNN architectures, including larger memory footprints, model sizes, higher FLOPs, and longer inference times [14]. Recent models employ billions of parameters to enhance feature representation. This often leads to overfitting and challenges in generalization, especially when the training data is limited or poorly structured. These issues become even more pronounced in cross-domain settings [15, 16] , where models must adapt to divergent data distributions.
@@
Image mixing based data augmentation techniques have proven effective for enhancing the generalization of deep learning models [17]. A wide array of such techniques has been introduced, including CutMix [18], Mixup [19], SnapMix [20], Co-Mixup [21], SaliencyMix [22], LGCOAMix [23], GuidedMixup [24], MixPro [25], TransMix [26], and PuzzleMix [19]. These methods typically involve combining portions of randomly selected images of corresponding labels through various mixing strategies to generate augmented samples - see top row in Fig. 1. By employing linear interpolation or other blending techniques, they create novel training images to enrich the dataset for addressing potential model overfitting [27, 28, 29, 30, 20, 31].
@@
Though effective, image mixing based data augmentation encounters challenges, such as the omission of salient regions and label ambiguities due to the random blending of image content [19]. To mitigate these issues, several approaches have incorporated saliency-driven methods, which prioritize the preservation of important regions by overlaying them onto less significant background areas of the target image [19, 22, 24, 32]. While these techniques offer improvements, they introduce considerable computational overhead and rely on the accuracy of the saliency methods, which is a limiting factor. Furthermore, by combining images from diverse categories, these techniques can still overlook crucial contextual information, leaving the problem of inadequate feature preservation unresolved.
@@
Figure 1: Representative examples of augmentation techniques. Top: Conventional image mixing methods like CutMix [27], Mixup [28], GridMix [33], SmoothMix [34], ResizeMix [35] and FMix [36] use interpolation or other strategies to mix source and target images, largely disregarding the natural appearance of the resulting image. Bottom: Our GenMix is a label-preserving technique like AugMix [37], PixMix [38] and IPMix [39]. However, it utilizes a generative diffusion model with custom tailored prompts for high quality augmentation. Three different outputs (1A, 2A and 3A) for the target are shown for GenMix.
@@
Diffusion Models (DMs) [40, 41, 42, 43, 44, 45] have recently emerged as powerful tools for image-to-image generation and editing. Some studies [46, 47, 48] have also explored using DM generated images to augment training data. However, we empirically observe limited performance gains from these approaches, with models trained on synthetic data occasionally underperforming compared to the baseline models without augmentation [47]. This can be attributed to the limited control over image generation of the DMs. Due to the model sensitivity to conditional prompts, generating complex scenes, layouts, and shapes becomes challenging [49, 50]. Poorly crafted prompts often yield unsuitable images for augmentation, which can lead to model adaption to an incorrect data distribution. This underscores the need for more precise prompt selection and better control mechanisms in the augmentation process to improve both domain generalization [51] and cross-domain performance.
@@
To address this, we propose GenMix, a generlizable data augmentation technique that leverages generative pre-trained diffusion models to compute diverse augmentation samples through carefully designed conditional prompts. GenMix creates hybrid images by blending original and generated content, maintaining essential semantics of the original image, see bottom row of Fig. 1. To further enrich structural diversity, we incorporate self-similarity fractals into the hybrid images, a strategy proven to enhance machine learning safety [38, 52]. This blending mitigates overfitting and improves model performance. Our strategy is unlike any previous attempts of directly using DM images for augmentation [47, 46]. It effectively incorporates a range of operations in its components to achieve outstanding eventual performance - see Table I. Our experiments show that GenMix surpasses current state-of-the-art (SOTA) augmentation methods by improving generalization and adversarial robustness. Additionally, it facilitates prompt-guided learning in both in-domain and cross-domain tasks, making it highly adaptable for diverse computer vision problems. Furthermore, our results confirm that GenMix is compatible with a wide range of datasets and it integrates seamlessly into existing SOTA frameworks.
@@
  • We explore the applicability of GenMix to both in-domain classification and domain adaptation tasks, introducing sets of bespoke conditional prompt for enhanced adaptability. Notably, GenMix is the first method to address both in-domain and domain adaptation tasks effectively.
  • We demonstrate with extensive experiments across eight datasets that GenMix outperforms existing SOTA augmentation methods in general classification, fine-grained classification, adversarial robustness, and cross-domain tasks.
@@
AutoMix
AdAutoMix
DiffuseMix
@@
Components
Source image
@@
Tasks
Adversarial robustness
@@
A preliminary version of this work was presented at a conference [53]. The present manuscript substantially extends that early effort by introducing several novel contributions. First, the DiffuseMix [48] lacks mechanisms for reliability and responsible image editing, we address this gap by proposing an unsupervised data augmentation strategy. Our method extracts robust features from generative images without requiring labels, thereby filtering out undesirable synthetic samples. Second, we introduce a novel label-preserving approach, termed PatchSwap, which exchanges patches between generative and natural images. Third, we design a smoother concatenation technique that fuses original and synthetic images through a smoothness factor, consistently improving model performance. Fourth, we develop new contextual prompts tailored for domain adaptation scenarios including open-partial, open-set, and partial-set adaptation. We demonstrate that training the state-of-the-art LEAD framework [54] with our proposed method yields remarkable gains across diverse categories. To the best of our knowledge, no existing approach jointly addresses both general image classification and domain adaptation tasks within a unified framework. Finally, we extend our empirical study to a wide range of architectures, spanning small-scale backbones (ResNet-18, ResNet-34, ResNet-50), large-scale backbones (ResNeXt-50, ResNet-101), and vision transformers (Swin Transformer, ConvNeXt), confirming the generality and scalability of our approach.
II Related Work
@@
Diffusion Models for Augmentation. Recently, there has been a growing interest in leveraging Diffusion Models (DMs) for data augmentation. Azizi et al. [47] proposed using text-to-image fine-tuned diffusion models to generate synthetic samples for ImageNet classification, demonstrating that incorporating these generated images into the training set can enhance classification performance. Similarly, Trabucco et al. [46] utilized off-the-shelf diffusion models to create diverse and semantically rich images via prompt engineering, aiming to improve image classification on various datasets. Furthermore, Li et al. [55] explored using diffusion models for data augmentation in conjunction with knowledge distillation, particularly in scenarios where real images are unavailable. These efforts underscore the potential of diffusion models to enrich data augmentation techniques.
Image Mixing for Augmentation. Image mixing has become a widely employed category of data augmentation techniques, enhancing the performance and robustness of deep learning models [56, 57, 58, 59]. Notable state-of-the-art methods in this category include CutMix, AugMix and PuzzleMix. CutMix [18] improves out-of-distribution generalization by randomly overlaying a patch from a source image onto a target image. AugMix [37] stochastically combines augmentation operations like equalize, posterize, and translate-x to introduce semantic diversity. PuzzleMix [19] refines mixup techniques by considering image saliency and local statistics. Similarly, another technique SaliencyMix [22] leverages saliency maps to focus on the most important image regions, preserving their integrity. Manifold Mixup [60] blends hidden network states during training, creating an interpolated state for improved feature representation. Another well-known method is PixMix [38] that extends AugMix [37] by mixing input images with fractal and feature visualization images, improving machine learning safety [61]. A comprehensive overview of various image-mixing methods, along with their components and task accomplishment is presented in Table I.
Automated Data Augmentation. Several researchers have also explored automated data augmentation for improving model performance. For instance, AutoAugment [62] utilizes reinforcement learning to discover optimal data augmentation strategies, while RandAugment [63] employs a set of randomized augmentation operations to enhance model generalization and robustness. AdaAug [64] has been introduced to efficiently learn adaptive augmentation policies that vary depending on the class or even the instance.
@@
Figure 2: Overview of the proposed filtration module without label supervision.
Figure 3: Overview of the proposed GenMix method. An input image and a randomly selected prompt are input to a pre-trained image editing diffusion model to obtain a generated image. Input and generated images are smoothly concatenated using a binary mask to obtain a hybrid image. Random fractal image is blended with this hybrid image to obtain the augmented image.
III Proposed Method
III-A Overview
Existing image mixing based data augmentation techniques may induce label ambiguities by pasting the source image contents onto the target image and consequently replacing either important regions of the foreground or the required background context [27, 24]. The central idea of our approach is to augment the source image only while preserving the key semantics and concatenating a portion of the source image with its generative counterpart. At the same time, our method introduces diverse object details and contexts in the resulting image for effective augmentation.
The schematics of the proposed GenMix technique is provided in Fig. 3. Our method comprises three main steps: prompt-guided image editing, seamless concatenation, selection of faithful images, and fractal interpolation. Initially, we use conditional prompts with a diffusion-based image editing module to compute generative counterparts of the source image. Then, a portion of the source image is concatenated with the rest of the generated image, forming a hybrid image. This approach ensures that the training model has access to the original visual content along with the generated one. Finally, the hybrid image is blended with a random fractal image, obtaining the final training image with a diverse structure. Improvements of our image by blending fractal is inspired by [38, 65]. However, in our case, the blending objective for augmentation is unique to our approach.
III-B GenMix Details
[…]
@@ III-B GenMix Details
Prompt Guided Image Editing: Our image editing step $\mathcal{E}(.)$ consists of a pre-trained image editing diffusion model that requires a prompt $p_{j}$ from a predefined set of $k$ prompts, $P=\{p_{1},p_{2},\dots,p_{k}\}$ where $j\in[1,k]$ along with the input training image $I_{i}$ and produces an edited counterpart image $\hat{I}_{ij}$. The image generation process in conventional diffusion models is often open-ended for image-to-image or text-to-image translation. On the other hand, image editing is guided by textual prompts. The aim is to obtain diverse image-to-image translations guided by textual prompts. In our case, as the goal is to achieve a slightly modified but not too different version of $I_{i}$, filter-like prompts are curated in $\mathcal{P}$ which do not alter the image drastically. Examples of the prompts used in GenMix are shown in Figure 4. The overall generation step can be represented as:
@@
Selection of Faithful Images: To ensure semantic consistency, we retain only those generated images whose similarity to their corresponding original image exceeds a dynamic threshold. Formally, the selection criterion for faithful images is defined as:
$\hat{I}_{ij}\in\mathcal{F}\quad\text{if}\quad\text{cos\_sim}\left(f_{m}(I_{i}),f_{m}(\hat{I}_{ij})\right)\geq\mu-2\sigma$ (3)
where $f_{m}(\cdot)$ is the DINOv2 feature extractor, $\mu$ and $\sigma$ denote the mean and standard deviation of pairwise cosine similarities computed among all original images $I_{i}$, and $\mathcal{F}$ denotes the set of faithful generated images retained for augmentation. As defined in Eq. (3), we discard dissimilar generations and retain only faithful images for downstream augmentation.
@@
TABLE II: Top-1 (%) performance of mixup methods on CIFAR-100, Tiny-ImageNet and ImageNet-1K. The previous SOTA miup methods results are taken from AdAutoMix [30] and reproduce DiffuseMix [53] results on new backbones.
CIFAR100 CIFAR100 Tiny-ImageNet ImageNet-1K
Method ResNet18 ResNeXt50 Swin-T ConvNeXt-T ResNet18 ResNeXt50 ResNet18 ResNet34 ResNet50
Vanilla [66] 78.04 81.09 78.41 78.70 61.68 65.04 70.04 73.85 76.83
MixUp [28] 79.12 82.10 76.78 81.13 63.86 66.36 69.98 73.97 77.12
CutMix [27] 78.17 81.67 80.64 82.46 65.53 66.47 68.95 73.58 77.17
SaliencyMix [22] 79.12 81.53 80.40 82.82 64.60 66.55 69.16 73.56 77.14
FMix [36] 79.69 81.90 80.72 81.79 63.47 65.08 69.96 74.08 77.19
PuzzleMix [19] 81.13 82.85 80.33 82.29 65.81 67.83 70.12 74.26 77.54
ResizeMix [35] 80.01 81.82 80.16 82.53 63.74 65.87 69.50 73.88 77.42
AutoMix [67] 82.04 83.64 82.67 83.30 67.33 70.72 70.50 74.52 77.91
AdAutoMix [30] 82.32 84.22 84.33 83.54 69.19 72.89 70.86 74.82 78.04
DiffuseMix [53] 82.78 84.88 85.26 84.38 70.47 73.62 71.62 75.23 78.64
GenMix 84.25 86.23 86.28 85.25 71.38 74.81 72.32 75.14 79.32
Algorithm 1 GenMix
[…]
@@
$A_{ijuv}=\lambda F_{v}+(1-\lambda)H_{iju},\text{ }0\leq\lambda<1$ (5)
where $\lambda$ is the interpolation factor. This results in the final augmented image $A_{ijuv}$ used to train or fine-tune deep neural networks. The overall augmentation process of GenMix can be represented as: